OpenStudy (aye_my_peeps):

WHAT IS 3x2 1/3? a. 7 b. 6 1/3 c. 6 d. 5 1/3

OpenStudy (aye_my_peeps):

can anyone help me plz

OpenStudy (caseysorensen):

yea

OpenStudy (anitas.):

i will help but do you have any guesses?

jabez177 (jabez177):

\[3\times2 \frac{ 1 }{ 3 }\]

OpenStudy (aye_my_peeps):

at first I thought it was 6 1/3

jabez177 (jabez177):

3 in fraction form is \[\frac{ 3 }{ 1 }\]

jabez177 (jabez177):

\[\frac{ 3 }{ 1 }\times2 \frac{ 1 }{ 3 }\] is your equation.

jabez177 (jabez177):

but we have to convert the mixed number.